OpenStudy (anonymous):

Divide. Write each quotient in simplest form. 2/3 ÷ 2 = _____ can somone please teach me how to do this!??

OpenStudy (anonymous):

when you divide a fraction you invert the number/fraction you are dividing by and then multiply think of 2 as 2/1 so we have 2/3 / 2/1 = 2/3 x 1/2 - inverting 2/1 = 2 / 6 = 1/3

OpenStudy (anonymous):

hmm okay so i divide 2/3 and then after that one i divide it again?

OpenStudy (anonymous):

Nope. 2/3 x 1/2, you multiply the numerators and denominators to get 2/6

OpenStudy (anonymous):

And 2/6 simplifies to 1/3

OpenStudy (anonymous):

but where did the 1/2 come from?

OpenStudy (anonymous):

When you divide, its the same as multiplying by a fraction.

OpenStudy (anonymous):

invert 2/1 to get 1/2 - 'turn upside down' eg divide by 3/4 = multiply by 4/3

OpenStudy (anonymous):

If I have 2/2 its the same as saying 2/1 x 1/2.

OpenStudy (anonymous):

would you like to try another one miss?

OpenStudy (anonymous):

im going to try to do thes on my own now but thx yu

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

ahh!! i got the hang of it! now!